OpenStudy (anonymous):

Use the following function. log4(x + 3) = 2 To solve the equation, first: Select one: a. Re-write the equation using the definition of a logarithm. b. Divide both sides by 2. c. Take the log of each side. d. None of the above.

OpenStudy (johnweldon1993):

I mean first I would rewrite it \[\large \frac{\log(x + 3)}{\log(4)} = 2\]

OpenStudy (solomonzelman):

Using the rule of a log in relation to an equation where a number is raised to an exponent of x. \(\normalsize\color{black}{ \log_AB=C~~~~~~~~~~~~~A^C=B}\)

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \log_4(x+3)=2}\) \(\normalsize\color{black}{ 4^2=(x+3)}\) \(\normalsize\color{black}{ 16=(x+3)}\) \(\normalsize\color{black}{ 13=x}\)

OpenStudy (solomonzelman):

I would be a bit more creative though, \(\normalsize\color{black}{ \log_4(x+3)=2}\) \(\normalsize\color{black}{ \log_4(x+3)=2~\log_44}\) \(\normalsize\color{black}{ \log_4(x+3)=\log_44^2}\) \(\normalsize\color{black}{ \log_4(x+3)=\log_416}\) \(\normalsize\color{black}{ \log_4\color{red}{(x+3)} =\log_4\color{red}{16} }\) and on ....

OpenStudy (solomonzelman):

:)